A bicycle tire has a pressure of 450. kPa at 20.0°C. Assuming the volume does not change, what is the new pressure at 35.0°C?
1
Expert's answer
2020-06-05T05:14:00-0400

P1 = 450. kPa

T1 = 20.0°C = 293 K

P2 = unknown

T2 = 35.0°C = 308 K

V = constant


Solution:

Since the volume does not change, Gay-Lussac's law can be used.

Gay-Lussac's gas law can be expressed as: P1/T1 = P2/T2

To find the new pressure, solve the equation for P2:

P2 = (P1 × T2) / T1

P2 = (450. kPa × 308 K) / (293 K) = 473. kPa

P2 = 473. kPa


Answer: 473 kPa is the new pressure at 35.0°C.